# On material-uniform elastic bodies with disclinations and their homogenization

### v2

@@ 3 Bodies with discrete defects
[A single disclination] A two-dimensional body with one disclination can be obtained by removing a sector from a two-dimensional annulus, and gluing the opposite edges of the sector. Start with a domain
@@
for some $r_{1}>r_{0}\geq 0$, and $\alpha\in(0,1)$. Denote $\beta=1-\alpha$. Endow this body with the trivial implant map $\mathcal{P}_{0}=\operatorname{Id}$. Now, glue the edges $\theta=0$ and $\theta=2\pi\alpha$ by considering the map
$\chi(r\cos\theta,r\sin\theta)=\left(r\cos\frac{\theta}{\beta},r\sin\frac{\theta}{\beta}\right).$
@@
$\begin{split}\mathcal{P}_{1}|_{\varphi=0^{+}}&=e_{1}\otimes dr+e_{2}\otimes\beta r\,d\varphi=e_{1}\otimes dx+\beta e_{2}\otimes dy=\left(\begin{matrix}1&0\\ 0&\beta\end{matrix}\right)\\ \mathcal{P}_{1}|_{\varphi=0^{-}}&=e_{1}\otimes\left(\cos(2\pi\beta)\,dr-\beta r\sin(2\pi\beta)\,d\varphi\right)+e_{2}\otimes\left(-\sin(2\pi\beta)\,dr+\beta r\cos(2\pi\beta)\,d\varphi\right)\\ &=\left(\begin{matrix}\cos(2\pi\beta)&\beta\sin(2\pi\beta)\\ -\sin(2\pi\beta)&\beta\cos(2\pi\beta)\end{matrix}\right),\end{split}$
@@
$\mathcal{P}_{1}|_{\varphi=0^{-}}\circ\mathcal{P}_{1}^{-1}|_{\varphi=0^{+}}=\left(\begin{matrix}\cos(2\pi\beta)&\sin(2\pi\beta)\\ -\sin(2\pi\beta)&\cos(2\pi\beta)\end{matrix}\right)=\left(\begin{matrix}\cos(2\pi\alpha)&-\sin(2\pi\alpha)\\ \sin(2\pi\alpha)&\cos(2\pi\alpha)\end{matrix}\right)\in\mathcal{G}.$

### v3

@@ 3 Bodies with discrete defects
Remark: There are cases which, according to Definition 3.1, the body contains no defects, yet, one might consider them as having defects: for example, if $\mathcal{M}\subset\mathbb{R}^{2}$ is an annulus, and $\mathcal{P}$ is a global reference map that is the derivative of a double cover map $f:\mathcal{M}\to\mathcal{M}$. Even though the holonomy in this case is trivial, one can think of this body as having a negative disclination of excess angle $2\pi$. In order to rule out these cases one should additionally require (at least in 2D) that for sub-body to contain no defects, each generator of the fundamental group of the body has to have turning number $1$ (see, e.g., [KMS15, Theorem 5], [KM23, Definition 3.1]). However, these subtleties are not essential to the rest of this paper, for which the simpler Definition 3.1 is sufficient.
@@
[A single disclination] A two-dimensional body with one disclination (of positive curvature) can be obtained by removing a sector from a two-dimensional annulus, and gluing the opposite edges of the sector. Start with a domain
@@
for some $r_{1}>r_{0}\geq 0$, and $\alpha\in(0,1)$. Endow this body with the trivial implant map $\mathcal{P}_{0}=\operatorname{Id}$. Now, glue the edges $\theta=0$ and $\theta=2\pi\alpha$ by considering the map
$\chi(r\cos\theta,r\sin\theta)=\left(r\cos\frac{\theta}{\alpha},r\sin\frac{\theta}{\alpha}\right).$
@@
$\begin{split}\mathcal{P}_{1}|_{\varphi=0^{+}}&=e_{1}\otimes dr+e_{2}\otimes\alpha r\,d\varphi=e_{1}\otimes dx+\alpha e_{2}\otimes dy=\left(\begin{matrix}1&0\\ 0&\alpha\end{matrix}\right)\\ \mathcal{P}_{1}|_{\varphi=0^{-}}&=e_{1}\otimes\left(\cos(2\pi\alpha)\,dr-\alpha r\sin(2\pi\alpha)\,d\varphi\right)+e_{2}\otimes\left(\sin(2\pi\alpha)\,dr+\alpha r\cos(2\pi\alpha)\,d\varphi\right)\\ &=\left(\begin{matrix}\cos(2\pi\alpha)&-\alpha\sin(2\pi\alpha)\\ \sin(2\pi\alpha)&\alpha\cos(2\pi\alpha)\end{matrix}\right),\end{split}$
@@
$\mathcal{P}_{1}|_{\varphi=0^{-}}\circ\mathcal{P}_{1}^{-1}|_{\varphi=0^{+}}=\left(\begin{matrix}\cos(2\pi\alpha)&-\sin(2\pi\alpha)\\ \sin(2\pi\alpha)&\cos(2\pi\alpha)\end{matrix}\right)\in\mathcal{G}.$